OpenStudy (anonymous):

The price of a suit decreased from $280 to $182 during a clearance sale. What is the percent of decrease in the price? Round to the nearest percent, if necessary. A. 54% B. 98% C. 65% D. 35%

OpenStudy (vera_ewing):

ok so it can't be B that's for sure...

OpenStudy (vera_ewing):

The price decreased by $98. The original price was $280 so it is the percentage decrease of this figure we are looking for. The calculation is:- 98/280 x 100 = 9800/280 = 35 so the decrease is 35%

OpenStudy (anonymous):

I think that it would be D.

OpenStudy (anonymous):

280-182=98 280*.35=98

OpenStudy (anonymous):

thnx everyone!!!

OpenStudy (vera_ewing):

no problem :) let me know if you have any more!